OpenStudy (anonymous):

The length of a rectangular box is 2 units less than its width. If the area of the box is 15 units square. Find the perimeter of the box.

OpenStudy (anonymous):

Perimeter of box? Length= w-2 Area= (w)(w-2)=15 \[w^2-2w=15\] \[w^2-2w-15=0\] w= -3, or 5 (but we can just throw out the -3 since the length and width can't be (-)) w=5, and l=3 Perimeter= 2(3)+2(5)=16 units